# Ambiguity-aware Truncated Flow Matching for Ambiguous Medical Image Segmentation

### v1

@@ Methods
Computing the intermediate prediction corresponding to timestep t (line 2 and 3 in Algorithm 1): The flow transformation follows an Optimal Transformation (OT) schedule (14), representing the shortest path between source and target distributions. Under the OT framework, the diffusion trajectory in the latent space forms a line segment. Therefore, we perform analytic geometry in the latent space: the intermediate state $X_{t}$ at timestep $t$ is computed by linear interpolation between the source endpoint $X_{T_{\text{trunc}}}$ and the target endpoint $X_{1}$. Then, using the direction vector of the segment $g_{t}(X_{t})$ and the position of $X_{t}$, a predicted result $x_{1t}$ is derived by projecting along the diffusion trajectory starting from $x_{t}$. (Further proof and analysis of line 3 in Algorithm 1 are provided in the appendix of extended version.)
@@
All training and inference procedures are conducted on a single RTX 3090 GPU with 24GB memory. SFM in ATFM is trained for 200 epochs on LIDC with GTR pretrained for 1000 epochs, and for 120 epochs on ISIC3 with a 400-epoch GTR. We set $\lambda=10^{-3}$ (i.e. $T=1000$) with a linear schedule for all experiments. Both GTR and ST-Net of the SFM are optimized using an Adam optimizer (28) with a learning rate of $10^{-4}$. Hyper-parameter $M$ is set to $20$ following (17) and $\alpha$ is set to $10^{-3}$ and $10^{-4}$ for LIDC and ISIC3 respectively according to hyper-parameter studies. (Detailed structures and implementation details of the proposed GTR and ST-Net are listed in the appendix of extended version.)
@@
Three metrics are utilized for comprehensive evaluation from three aspects: For segmentation distribution, we utilize the Generalised Energy Distance (GED) (18) to evaluate the alignment among distribution of predictions and ground-truths. For sample fidelity, we utilize the Hungarian-Matching Intersection-over-Union (HM-IoU) (20) to provide an accurate representation of the performance on segmentation across all predictions. For individual segmentation accuracy, we utilize the Maximum Dice Matching (MDM) (1) to evaluate the best Dice scores between each prediction result and each ground truth. We denote the metrics with subscript n to represent the metrics calculated using n samples. Results are reported as mean ± standard deviation over five independent runs. (Detailed description and implementation details of the evaluation metrics are listed in the appendix of extended version.)
@@
Fig. 4 and Fig. 5 further compare visual results among advanced methods and proposed ATFM for LIDC dataset and ISIC3 dataset, respectively. Predictions from ATFM more faithfully reflect the range of plausible annotations and better preserve fine-grained structures, indicating its superior alignment with ambiguous and detailed ground truths, which is an essential and necessary requirement in AMIS tasks. (More qualitative results of both datasets are illustrated in the appendix of extended version.)
@@
Table 4 and Fig. 6(a) show the conducted ablation study on both datasets evaluating five model variants: GTR with activation layers (Act. GTR), SFM (with and without $L_{\text{SF}}$), and ATFM (with and without $L_{\text{SF}}$). ATFM outperformed both Act. GTR and SFM by a minimum of $10\%$ and $6\%$ on both metrics, highlighting the benefit of Data-Hierarchical Inference and the high effectiveness and fidelity of proposed ATFM provided by GTR. The performance gap of an average of $11\%$ between models with and without $L_{\text{SF}}$ emphasizes the importance of semantics consistency modeling, underscoring the role of $L_{\text{FM}}$ and SFM in preserving plausibility when enhancing diversity.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62501195, 62272135 and 62372135, and the Key Research & Development Program of Heilongjiang Province under Grants 2023ZX01A08 and 2024ZX12C23, and the Natural Science Foundation of Heilongjiang Province under Grant LH2024F019.

### v2

@@ Methods
Computing the intermediate prediction corresponding to timestep t (line 2 and 3 in Algorithm 1): The flow transformation follows an Optimal Transformation (OT) schedule (14), representing the shortest path between source and target distributions. Under the OT framework, the diffusion trajectory in the latent space forms a line segment. Therefore, we perform analytic geometry in the latent space: the intermediate state $X_{t}$ at timestep $t$ is computed by linear interpolation between the source endpoint $X_{T_{\text{trunc}}}$ and the target endpoint $X_{1}$. Then, using the direction vector of the segment $g_{t}(X_{t})$ and the position of $X_{t}$, a predicted result $x_{1t}$ is derived by projecting along the diffusion trajectory starting from $x_{t}$.
@@
All training and inference procedures are conducted on a single RTX 3090 GPU with 24GB memory. SFM in ATFM is trained for 200 epochs on LIDC with GTR pretrained for 1000 epochs, and for 120 epochs on ISIC3 with a 400-epoch GTR. We set $\lambda=10^{-3}$ (i.e. $T=1000$) with a linear schedule for all experiments. Both GTR and ST-Net of the SFM are optimized using an Adam optimizer (28) with a learning rate of $10^{-4}$. Hyper-parameter $M$ is set to $20$ following (17) and $\alpha$ is set to $10^{-3}$ and $10^{-4}$ for LIDC and ISIC3 respectively according to hyper-parameter studies.
@@
Three metrics are utilized for comprehensive evaluation from three aspects: For segmentation distribution, we utilize the Generalised Energy Distance (GED) (18) to evaluate the alignment among distribution of predictions and ground-truths. For sample fidelity, we utilize the Hungarian-Matching Intersection-over-Union (HM-IoU) (20) to provide an accurate representation of the performance on segmentation across all predictions. For individual segmentation accuracy, we utilize the Maximum Dice Matching (MDM) (1) to evaluate the best Dice scores between each prediction result and each ground truth. We denote the metrics with subscript n to represent the metrics calculated using n samples. Results are reported as mean ± standard deviation over five independent runs.
@@
Fig. 4 and Fig. 5 further compare visual results among advanced methods and proposed ATFM for LIDC dataset and ISIC3 dataset, respectively. Predictions from ATFM more faithfully reflect the range of plausible annotations and better preserve fine-grained structures, indicating its superior alignment with ambiguous and detailed ground truths, which is an essential and necessary requirement in AMIS tasks.
@@
Table 4 and Fig. 6(a) show the conducted ablation study on both datasets evaluating five model variants: GTR with activation layers (A), SFM with and without $L_{\text{SF}}$ (B and C), and ATFM with and without $L_{\text{SF}}$ (D and E). ATFM outperformed both Act. GTR and SFM by a minimum of $10\%$ and $6\%$ on both metrics, highlighting the benefit of Data-Hierarchical Inference and the high effectiveness and fidelity of proposed ATFM provided by GTR. The performance gap of an average of $11\%$ between models with and without $L_{\text{SF}}$ emphasizes the importance of semantics consistency modeling, underscoring the role of $L_{\text{FM}}$ and SFM in preserving plausibility when enhancing diversity.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62501195, 62272135, 62372135 and 82527807, and the Key Research & Development Program of Heilongjiang Province under Grants 2023ZX01A08 and 2024ZX12C23, and the Natural Science Foundation of Heilongjiang Province under Grant LH2024F019.